Two capacitors connected in parallel produce an equivalent capacitance of 40.0 μF but when connected in series the equivalent capacitance is only 7.1 μF .

What is the individual capacitance of each capacitor?
Answer #1

Let C_1\ and\ C_2 be capacitances of the two capacitors. When connected in parallel, the equivalent capacitance is given by

C_1+C_2=40\ \mu F\ \ \ \ \ \ \ \ (1)

When connected in series the equivalent capacitance is given by

\frac{C_1C_2}{C_1+C_2}=7.1\ \mu F

Using equation (1) in the denominator of LHS

\frac{C_1C_2}{40\ \mu F}=7.1\ \mu F

C_1C_2=40\times7.1

Again using equation (1) we get C_1=40-C_2 ,

(40-C_2)C_2=284

40C_2-C_2^2=284

C_2^2-40C_2+284=0

This is a quadratic equation in C2, the roots of a quadratic equation ax^2+bx+c=0 are

x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}

The roots of the above quadratic equation are

C_2=\frac{-(-40)\pm \sqrt{(-40)^2-4\times 1\times 284}}{2\times 1}

C_2=30.77\ \mu F,\ 9.23\ \mu F

Using (1) the other capacitance for each case is

C_1= 9.23\ \mu F,\ 30.77\ \mu F

Therefore, the two capacitors are 9.23\ \mu F,\ 30.77\ \mu F .